# PoTable: Programming Standardly on Table-based Reasoning Like a Human Analyst

### v1

@@ 1 Introduction
Tables are widely applied in various scenarios (e.g., healthcare (Ghasemi & Amyot, 2016), finance (Li et al., 2021)), since they can visually present the core information in various types of scientific documents (e.g., articles, reports, websites) (Embley et al., 2006) through a structured format. With the growing development of AI techniques, there has been an increasing demand for automated table processing, attracting significant attention from both academia and industry (Borisov et al., 2022). Recently, the evolution of Large Language Model (LLM) (Zhao et al., 2023) has raised a brand new prompting paradigm for table processing (Lu et al., 2024). This training-free method facilitates complex understanding and reasoning procedures in table question answering (Pasupat & Liang, 2015), table fact verification (Chen et al., 2020) and other downstream tasks (shown in Figure 1(a)).
@@
Throughout the history of humankind, tools have been regarded as the crystallization of human wisdom and a core factor in social productivity development (Washburn, 1960). This consensus has inspired LLM-based techniques to go a step further in simulating more extensive human behavior, i.e., collaborating with symbolic tools to overcome LLMs’ inherent limitations (Qu et al., 2024). In table processing, two unique challenges have been issued in earlier studies (Lu et al., 2024; Dong & Wang, 2024): (1) Tables are structured in two-dimension, leading to unstable memorization of LLMs trained in next-token prediction mode (Sui et al., 2024). (2) Table-based reasoning inevitably involves logical and arithmetic operations, and LLMs may produce misleading results due to their limited calculation abilities. Nevertheless, with a rising trend to utilize databases (Li et al., 2023b), Python (Chen et al., 2022; Gao et al., 2023) and other symbolic tools as assistants, recent approaches effectively reduce table processing errors and misleading computational results by storing the tabular data into internal structure types (e.g., arrays, database tables) and executing syntactic computation commands (e.g., SQL, Python code) (Cheng et al., 2023; Cao et al., 2023; Nahid & Rafiei, 2024).
@@
Despite promising results of the existing studies, there is still room for improvement in simulating human cognitive behavior in table-based reasoning. As shown in Figure 1(b), the human analyst follows a standard tabular data analysis process, considering all possible open-world operations in each stage to reach the local target and finally solve the given tabular task. In general, all the above stages and operations are implemented through a symbolic tool, such as Python with pandas [^1] https://pandas.pydata.org/ package. However, existing studies ignore the standardization of logical split and only consider a few specialized operations for execution. Specifically, some studies attempt to generate the overall program in one go for execution (Cheng et al., 2023), while a few studies iteratively make decisions from pre-defined operation pools without logical boundary (Wang et al., 2024b). On the other hand, partial studies generate executable SQL or Python code of only sub-table and cell content extraction (Ye et al., 2023; Nahid & Rafiei, 2024), while a few studies only specify several commonly used operations for execution (Wang et al., 2024b). In summary, existing executor-based approaches still suffer from non-standard logical splits and constrained operation pools in tool execution. More effective utilization of symbolic tools is required to achieve better performance.
@@
It can be clearly seen that PoTable follows a human-like analysis stage split and extends the constrained operation pool into an open-world space for symbolic tool execution. As a result, PoTable produces superior reasoning results with highly accurate, steply commented and completely executable code, which demonstrates its effectiveness and explainability. Finally, we conduct extensive experiments over three evaluation datasets from two public benchmarks of table-based reasoning tasks on two backbones. PoTable achieves much more outstanding accuracy results than all LLM-based baselines. In particular, GPT-based PoTable achieves over 4% higher absolute accuracy than runner-ups on all evaluation datasets. All experimental results and analyses validate the strong effectiveness of PoTable. In summary, our main contributions can be listed as follows:
  • We propose PoTable, a novel table-based reasoning method that programs standardly to simulate a human analyst. Our method follows a human-like analysis stage split and extends the constrained operation pool into an open-world space for symbolic tool execution.
  • We integrate a Python interpreter as the real-time executor along with an LLM-based operation planner and code generation. PoTable produces high-quality reasoning results with highly accurate, steply commented and completely executable programs, demonstrating its strong effectiveness and explainability.
  • Experimental results over three evaluation datasets from two public benchmarks of table-based reasoning tasks show the outstanding performance of PoTable.
@@ 3.1 Experimental Setup
Baselines. We select four competitive LLM-based approaches as baselines for comparison. Binder (Cheng et al., 2023) is a neural-symbolic framework that maps the reasoning task into a specific program and then executes the program binding LLM as a unified API to extend its grammar coverage and tackle the commands that cannot be executed normally. Dater (Ye et al., 2023) first decomposes the table into sub-evidence with column and row selection through LLM queries and then decomposes the question into simpler sub-questions through intermediate SQL generation, followed by a joint reasoning stage with simplified tables and questions. Chain-of-Table (Wang et al., 2024b) pre-defines several common atomic operations for dynamic selection by the LLM, forming an operation chain to process the table with pre-defined code to simplify the table for the final LLM answer querying. TabSQLify (Nahid & Rafiei, 2024) leverages Text-to-SQL to decompose the table into sub-tables and conduct comprehensive reasoning and answer generation through LLM queries. All these selected approaches are competitive as LLM-based baselines on table-based reasoning.
@@ 4 Related Work
Table Processing with Language Models. Table processing has been a popular research domain over the past decade. Before the era of LLMs, numerous efforts were made to process tables with pre-trained language models. TaPas (Herzig et al., 2020) extends BERT (Devlin et al., 2019) by conducting masked pre-training with joint encoding of questions and flattening tables. TaBERT (Yin et al., 2020) combines content snapshot and vertical attention based on BERT to obtain joint textual and tabular representations for further understanding. TUTA (Wang et al., 2021) enhances transformers (Vaswani et al., 2017) with structure-aware mechanisms to effectively capture spatial, hierarchical and semantic information. TAPEX (Liu et al., 2022) pre-trains BART (Lewis et al., 2020) on a large synthetic SQL dataset to imitate the SQL executor that better understands tabular structure information. With the development of LLMs, the paradigm of table processing has been deeply revolutionized, especially in tabular data encoding and reasoning. In prompting methods, Sui et al. (2024) designs a benchmark to evaluate the structural understanding capabilities of LLMs, followed by a novel self-augmentation for effective structural prompting. Dater (Ye et al., 2023) and DIN-SQL (Pourreza & Rafiei, 2023) adopt task decomposition for better understanding with simplified queries, while Chain-of-Table (Wang et al., 2024b) defines atomic operations for dynamic selection in CoT prompting. Some other approaches explore training or tuning LLMs as generalists. TableLlama (Zhang et al., 2024) develops an open-source tabular LLM by fine-tuning Llama 2-7B (Touvron et al., 2023) with LongLoRA (Chen et al., 2024b), while Table-LLAVA (Zheng et al., 2024) trains a multi-modal tabular LLM that can handle table images as vision inputs.
@@
Table Processing with Symbolic Tools. Symbolic tools have been widely utilized as assistants to produce more accurate and robust mid-results in LLM-based table reasoning scenarios. Most studies adopt databases and Python as affiliated executors to interact with LLMs. Binder (Cheng et al., 2023) and Cao et al. (2023) parse the tasks into integral SQL or Python programs for further execution, incorporating LLM-assistant APIs to handle abstract code blocks for complete execution. TabSQLify (Nahid & Rafiei, 2024) generates SQL queries to extract sub-tables and executes them to get simplified tables for further LLM reasoning. Some works target boosting the code generation ability for tabular reasoning and other scenarios. TroVE (Wang et al., 2024a) asks the code LLMs to curate reusable high-level functions and use them to write solutions for Python execution on the table question answering and other tasks, while Self-Debugging (Chen et al., 2024a) teaches LLMs to debug their predicted SQL or Python programs on Text-to-SQL (Yu et al., 2018) and other tasks. Recently, research in LLM-based table reasoning has been extended into more sophisticated tools environments and more advanced reasoning tasks. SheetCopilot (Li et al., 2023a) and SpreadsheetBench (Ma et al., 2024) address a novel spreadsheet manipulation task, which maneuvers table analysis software like Microsoft Excel [^4] https://www.microsoft.com/zh-cn/microsoft-365/excel to generate step-by-step solutions for simulated execution. MatPlotAgent (Yang et al., 2024) addresses the task of scientific data visualization, which includes a code agent integrating Matplotlib [^5] https://matplotlib.org/ responsible for generating the code to plot figures from input tables. As a result, symbolic tool utilization has become a crucial component in table processing.

### v2

@@ 1 Introduction
Tables are widely applied in various scenarios (e.g., healthcare (Ghasemi & Amyot, 2016), finance (Li et al., 2021)). They can visually present the core information in various types of scientific documents (e.g., articles, reports, websites) (Embley et al., 2006) through a structured format, benefiting downstream tasks (e.g., time-series forecasting (Cheng et al., 2023a)). With the growing development of AI techniques, there has been an increasing demand for automated table processing, attracting significant attention from both academia and industry (Borisov et al., 2022). Recently, the evolution of Large Language Model (LLM) (Zhao et al., 2023; Cheng et al., 2024) has raised a brand new table processing strategy (Lu et al., 2024). The prompting method facilitates complex understanding and reasoning procedures in table question answering (Pasupat & Liang, 2015), table fact verification (Chen et al., 2020) and other downstream tasks (shown in Figure 1(a)).
@@
Throughout the history of humankind, tools have been regarded as the crystallization of human wisdom and a core factor in social productivity development (Washburn, 1960). This consensus has inspired LLM-based techniques to go a step further in simulating more extensive human behavior, i.e., collaborating with symbolic tools to overcome LLMs’ inherent limitations (Qu et al., 2024). In table processing, two unique challenges have been issued in earlier studies (Lu et al., 2024; Dong & Wang, 2024): (1) Tables are structured in two-dimension, leading to unstable memorization of LLMs trained in next-token prediction mode (Sui et al., 2024). (2) Table-based reasoning inevitably involves logical and arithmetic operations, and LLMs may produce misleading results due to their limited calculation abilities. Nevertheless, with a rising trend to utilize databases (Li et al., 2023b), Python (Chen et al., 2022; Gao et al., 2023) and other symbolic tools as assistants, recent approaches effectively reduce table processing errors and misleading computational results by storing the tabular data into internal structure types (e.g., arrays, database tables) and executing syntactic computation commands (e.g., SQL, Python code) (Cheng et al., 2023b; Cao et al., 2023; Nahid & Rafiei, 2024).
@@
Despite promising results of the existing studies, there is still room for improvement in simulating human cognitive behavior in table-based reasoning. As shown in Figure 1(b), the human analyst follows a standard tabular data analysis process, considering all possible open-world operations in each stage to reach the local target and finally solve the given tabular task. In general, all the above stages and operations are implemented through a symbolic tool, such as Python with pandas [^1] https://pandas.pydata.org/ package. However, existing studies ignore the standardization of logical split and only consider a few specialized operations for execution. Specifically, some studies attempt to generate the overall program in one go for execution (Cheng et al., 2023b), while a few studies iteratively make decisions from pre-defined operation pools without logical boundary (Wang et al., 2024c). On the other hand, partial studies generate executable SQL or Python code of only sub-table and cell content extraction (Ye et al., 2023; Nahid & Rafiei, 2024), while a few studies only specify several commonly used operations for execution (Wang et al., 2024c). In summary, existing executor-based approaches still suffer from non-standard logical splits and constrained operation pools in tool execution. More effective utilization of symbolic tools is required to achieve better performance.
@@
It can be clearly seen that PoTable follows a human-like analysis stage split and extends the constrained operation pool into an open-world space for symbolic tool execution. As a result, PoTable produces superior reasoning results with highly accurate, steply commented and completely executable code, which demonstrates its effectiveness and explainability. Finally, we conduct extensive experiments over three evaluation datasets from two public benchmarks of table-based reasoning tasks on two backbones. PoTable achieves much more outstanding accuracy results than all LLM-based baselines. In particular, GPT-based PoTable achieves over 4% higher absolute accuracy than runner-ups on all evaluation datasets.
Our main contributions can summarized as follows: (1) We propose PoTable, a novel table-based reasoning method that programs standardly to simulate a human analyst. Our method follows a human-like analysis stage split and extends the constrained operation pool into an open-world space for symbolic tool execution. (2) PoTable produces high-quality reasoning results with highly accurate, steply commented and completely executable programs by integrating an LLM and a real-time Python interpreter, demonstrating its strong effectiveness and explainability. (3) Experimental results over three evaluation datasets from two public benchmarks of table-based reasoning tasks show the outstanding performance of PoTable.
@@ 3.1 Experimental Setup
Baselines. We select four competitive LLM-based approaches as baselines for comparison. Binder (Cheng et al., 2023b) is a neural-symbolic framework that maps the reasoning task into a specific program and then executes the program binding LLM as a unified API to extend its grammar coverage and tackle the commands that cannot be executed normally. Dater (Ye et al., 2023) first decomposes the table into sub-evidence with column and row selection through LLM queries and then decomposes the question into simpler sub-questions through intermediate SQL generation, followed by a joint reasoning stage with simplified tables and questions. Chain-of-Table (Wang et al., 2024c) pre-defines several common atomic operations for dynamic selection by the LLM, forming an operation chain to process the table with pre-defined code to simplify the table for the final LLM answer querying. TabSQLify (Nahid & Rafiei, 2024) leverages Text-to-SQL to decompose the table into sub-tables and conduct comprehensive reasoning and answer generation through LLM queries. All these selected approaches are competitive as LLM-based baselines on table-based reasoning.
@@ 4 Related Work
Table Processing with Language Models. Table processing has been a popular research domain over the past decade. Before the era of LLMs, numerous efforts were made to process tables with pre-trained language models. TaPas (Herzig et al., 2020) extends BERT (Devlin et al., 2019) by conducting masked pre-training with joint encoding of questions and flattening tables. TaBERT (Yin et al., 2020) combines content snapshot and vertical attention based on BERT to obtain joint textual and tabular representations for further understanding. TUTA (Wang et al., 2021) enhances transformers (Vaswani et al., 2017) with structure-aware mechanisms to effectively capture spatial, hierarchical and semantic information. TAPEX (Liu et al., 2022) pre-trains BART (Lewis et al., 2020) on a large synthetic SQL dataset to imitate the SQL executor that better understands tabular structure information. With the development of LLMs, the paradigm of table processing has been deeply revolutionized, especially in tabular data encoding and reasoning. In prompting methods, Sui et al. (2024) designs a benchmark to evaluate the structural understanding capabilities of LLMs, followed by a novel self-augmentation for effective structural prompting. Dater (Ye et al., 2023) and DIN-SQL (Pourreza & Rafiei, 2023) adopt task decomposition for better understanding with simplified queries, while Chain-of-Table (Wang et al., 2024c) defines atomic operations for dynamic selection in CoT prompting. TableTime (Wang et al., 2024a) holds perspectives of table encoding to address the time-series classification problem. Some other approaches explore training or tuning LLMs as generalists. TableLlama (Zhang et al., 2024) develops an open-source tabular LLM by fine-tuning Llama 2-7B (Touvron et al., 2023) with LongLoRA (Chen et al., 2024b), while Table-LLAVA (Zheng et al., 2024) trains a multi-modal LLM that receives table images as vision inputs.
@@
Table Processing with Symbolic Tools. Symbolic tools have been widely utilized as assistants to produce more accurate and robust mid-results in LLM-based table reasoning scenarios. Most studies adopt databases and Python as affiliated executors to interact with LLMs. Binder (Cheng et al., 2023b) and Cao et al. (2023) parse the tasks into integral SQL or Python programs for further execution, incorporating LLM-assistant APIs to handle abstract code blocks for complete execution. TabSQLify (Nahid & Rafiei, 2024) generates SQL queries to extract sub-tables and executes them to get simplified tables for further LLM reasoning. Some works target boosting the code generation ability for tabular reasoning and other scenarios. TroVE (Wang et al., 2024b) asks the code LLMs to curate reusable high-level functions and use them to write solutions for Python execution on the table question answering and other tasks, while Self-Debugging (Chen et al., 2024a) teaches LLMs to debug their predicted SQL or Python programs on Text-to-SQL (Yu et al., 2018) and other tasks. Recently, research in LLM-based table reasoning has been extended into more sophisticated tools environments and more advanced reasoning tasks. SheetCopilot (Li et al., 2023a) and SpreadsheetBench (Ma et al., 2024) address a novel spreadsheet manipulation task, which maneuvers table analysis software like Microsoft Excel [^4] https://www.microsoft.com/zh-cn/microsoft-365/excel to generate step-by-step solutions for simulated execution. As a result, symbolic tool utilization has become a crucial component in table processing.